In a family there are two cars. In a given week, the first car gets an average of 40 miles per gallon, and the second car gets 35 miles per gallon. The two cars combined drive a total of 1650 miles in that week, for a total gas consumption of 45 gallons. How many gallons were consumed by each of the two cars that week?




Answer by checkley71(8403)

x/40+(1650-x)/35=45 or [(35x)+40(1650-x)]/40*35=45 or (35x+66000-40x)/1400=45 or (-5x+66000)/1400=45 or -5x+66000=63000 or -5x=-3000 or x=3000/5 or x=600 so
600 miles driven @40 mpg & 1650-600 or 1050 miles @ 35mpg then
600/40+1050/35=45 or 15+30=45 gallons